OpenStudy (anonymous):

Hi I need some help please. If an object is given an initial velocity of v0 feet per second from a height of s0 feet, then it's height s after t seconds is given by the formula S= -16t^(2) + v[0]t + s[0]. If a pine cone falls from the top of a 112 foot tree then how many seconds does it take to reach the earth? Round the answer to one decimal place

OpenStudy (amistre64):

initial velocity is 0, unless that was a powerful squirrel.... and initial height is 112; just gotta plug and chug it

OpenStudy (amistre64):

112.... 1,112; 4, 28,..4.4.7 4 sqrt(7) ------- = sqrt(7)/8 right? 32

OpenStudy (amistre64):

0.3 secs i think

OpenStudy (amistre64):

that doesnt look right :)

OpenStudy (amistre64):

16t^2 - 112 16(t^2-7)=0 t = sqrt(7) then.... lets try it :) -16(7) +112 = 0

OpenStudy (amistre64):

thats better

OpenStudy (amistre64):

2.64...

OpenStudy (anonymous):

Thanks

OpenStudy (angela210793):

This is physics....(and math) i don't like it....

OpenStudy (amistre64):

without physics; where would we be ;) (olivia newton john)

OpenStudy (angela210793):

I'd b still in Albania i know this 4 sure :P:P:P...

OpenStudy (amistre64):

Albania eh... is that were albinos come from? no offense

OpenStudy (angela210793):

Huh? nooo..Albinos is deficiency of vitamin D(if i'm not wrong) Albania is in EUROPE!!! (south-east) close to Greece :P:P:P:P:P

OpenStudy (angela210793):

While i was eating dinner i was thinking abt albinos and i understood that i have made a terrible mistake and i am sorry....:(:(:( that is caused because of deficiency of melanin not vit.D :)